a large metal cylinder of volume V contains helium at pressure 20 atmospheres (atm) or about 2.026 x 10(5)Pa when its temperature is 290 Kelvins about room temperature). What is the pressure of the helium if the temperature increases to 310 kelvins?

a. 21.2
b. 21.3
c. 21.4
d. 21.5
e. 21.6
Answer #1

If volume is constant, and initial pressure and temperatures are P1 and T1 respectively,while final pressure and temperatures are P2 and T2 respectively, then P1/P2=T1/T2

Here,P1=20 atm, T1=290 K, T2=310 K.

So, 20/P2=290/310, where P2 is the pressure at T2

=>P2=20*310/290=21.38 atm= 21.4 atm (rounded off to 1 decimal place)

So,required pressure is 21.4 atm.

So,option c is correct.